New Spell

1 Detection Spell

Beispielzauber

Dies ist ein virtueller Zauber, der eigentlich nur schwachsinnig ist. Zweite Zeile des Beispieles

Type:Physical  Range:Limited  Target:6
Damage:—  Duration:Sustained  Drain:(K/2)+1 M